OpenStudy (anonymous):

what is 23/20 as a percent

OpenStudy (anonymous):

115%

OpenStudy (anonymous):

can u please explain?:)

OpenStudy (saifoo.khan):

Sure, \[\frac{23}{20} \times 100\]

OpenStudy (anonymous):

oh ok that makes sense thanks

OpenStudy (saifoo.khan):

Same as with last one, \[\frac{1}{200} \times 100\]

OpenStudy (anonymous):

\[23/20=1.15\times100=115 perent\]